Let f(x)=3x-5. Find
(A) f(3) (B) f(h)
(C) f(3)+f(h) (D) f(3+h)
Please help! Thank you in advance!

Answer by Edwin McCravy(20086)

Let f(x)=3x-5. Find
(A) f(3)
Substitute 3 for x and simplify

    f(x)=3x-5
    f(3)=3(3)-5
    f(3)=9-5
    f(3)=4 

(B) f(h)
Substitute h for x

    f(x)=3x-5
    f(h)=3h-5    

(C) f(3)+f(h)
Add the answers to (A) and (B) and simplify:

    f(3)+f(h) = 4+3h-5
    f(3)+f(h) = -1+3h

(D) f(3+h)
Substitute 3+h for x and simplify

    f(x)=3x-5
    f(3+h)=3(3+h)-5
    f(3+h)=9+3h-5
    f(3+h)=4+3h

Edwin